Marie mixes 16 liters of 78% acid solution with a 25% acid solution to make a 41% acid solution.
How many liters of 25% solution did she use?



8 liters


12 liters


13.7 liters


37 liters


Answer by TimothyLamb(4379)

x = liters of 78%
y = liters of 25%
---
x + y = 16
78x + 25y = 41*16
---
put the system of linear equations into standard form
---
x + y = 16
78x + 25y = 656
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = liters of 78% = 4.8301887
y = liters of 25% = 11.169811
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---